# Box-Free Model Watermarks Are Prone to Black-Box Removal Attacks

### v2

@@ 1 Introduction
Common goals of DNN intellectual property protection are ownership verification and model stealing tracking (also known as model extraction or surrogate attack). Both goals can be effectively achieved through DNN watermarking [5, 6, 7]. Existing DNN watermarking methods can be classified into white-box, black-box, and box-free methods [5]. In white-box methods, the watermark is embedded directly into the model’s parameters, which requires access to the model’s internal structure for watermark extraction. While this approach allows for accurate watermark extraction due to direct access to the model parameters, it sacrifices the model privacy. Meanwhile, providing access to a model’s structure and parameters for extraction may also violate security protocols, leading to the potential alteration or removal of the watermark. In contrast, black-box methods circumvent the need to access the model’s internal information. Instead, they employ specific inputs, referred to as triggers, which can alter the model’s behavior. This is also termed as backdoor watermarking, and the verification is done by analyzing model inputs and outputs.
@@
The existing white- and black-box methods are primarily employed for the protection of classification models. We note that classification models only generate low entropic outputs, e.g., the categorical confidence scores, or even merely the top-1 label. This aligns with the essence of white- and black-box methods that embed watermarks into model parameters and behaviors, respectively. In contrast, for generative models, e.g., image-to-image models, the outputs contain high entropic image content, allowing the models to generate watermarked outputs, known as box-free watermarking [8]. In this type of methods, a watermark hiding network and an extraction network are respectively designed and can be trained jointly with the original model. Then, the protected model (original model plus hiding network) generates processed and watermarked images, while the extraction network, when queried by (stolen) model-generated images, yields the watermark, or otherwise yields a noisy or non-informative pattern. Box-free methods excel at their flexibility as they do not require access to model parameters. The output images are also called containers [8].
@@
The existing works on box-free methods have focused on watermark fidelity, robustness against surrogate attacks, robustness against normal image processing, and extension to different generative tasks [9, 10, 11]. Despite the current research progress, in this paper, we make the first attempt to reveal an overlooked vulnerability of box-free methods against black-box watermark removal attacks. Focusing on deep generative models for low-level image processing tasks, we consider three scenarios. Scenario 1: The victim model, consisting of an image processing network (IPNet), a watermark hiding network (HNet), and a watermark extraction network (ENet), is secured in a black box, whose API are available for query, and ENet only uses ReLU as activation. We then propose an extractor-gradient-guided (EGG) watermark removal attack and show its effectiveness. Scenario 2: It reflects a more challenging yet practical situation which is almost the same as Scenario 1 except that ENet uses arbitrary and unknown activation functions. In this scenario, ENet’s gradients can be estimated via querying, according to studies in adversarial attacks [^1] In black-box adversarial attack literature [12, 13, 14], the gradient information of the black-box victim model is estimated via querying. While this is beyond the scope of this paper, it validates our Scenario 2 in which the gradients of ENet can be reverse-engineered., and we develop an EGG remover under this condition. Scenario 3: It is the most challenging scenario in which the victim model is in a black box while ENet API is not provided. Since the original functionality of IPNet is known, we resort to a transferability-based attack, which trains a private set of proxy models denoted by $\text{HNet}^{\prime}$, and $\text{ENet}^{\prime}$, respectively. Then, we show that the remover trained for the proxy models can be transferred to attack the victim model.
@@
In the above three scenarios, the proposed EGG removers, as well as the transferable remover, when fed with container images, can effectively remove the embedded watermarks while preserving the image quality. We note that the proposed removers are also image-to-image models, which turn a container image into a watermark-free one. This differs from existing image-processing-based removal (e.g., adding noise, performing transformations, etc., usually ineffective) [15] and white-box ones (e.g., fine-tuning and compression, less practical than black-box ones) [16]. The contributions of this paper are summarized as follow:
@@
  • We reveal the vulnerabilities of the existing box-free model watermarking method against black-box removal attacks. This takes a distinctive perspective compared to most existing works dealing only with white-box removal such as fine-tuning and compression. This indicates that even black-box protection cannot effectively prevent watermark removal.
@@ 2.2 Black-Box Watermarking
Black-box model watermarking is also known as backdoor watermarking, whose watermark extraction process only requires the observation of the trigger-output relationships. Most existing backdoor watermarks are in the form of special trigger-label mappings, designed specially for deep classification models. The triggers can be both in- and out-of-distribution samples [18, 23, 24, 25, 26], while the trigger labels can be drawn from either the original categories or a novel category [27, 28]. However, Huang et al. [11] have pointed out the limitations of existing black-box methods in protecting generative models, noting that these methods are not applicable to noise-to-image models. In addition, black-box methods are unable to protect the generated images, which are also considered intellectual properties. In summary, while black-box watermarking could be more secure than white-box watermarking, it is unsuitable for safeguarding generative models. We note that Quan et al. [29] have made the first attempt to backdoor watermark generative models.
@@ 2.3 Box-Free Watermarking
Box-free watermarking is specifically designed for generative models. It embeds invisible watermarks into the generated outputs, and both the model and the outputs can be protected. In this line of research, Yu et al. [30] proposed an indirect approach to protect image-to-image generative adversarial networks (GANs). It first watermarks the training data and then, based on transferability, extracts the watermarks from images generated by the trained GANs using a pretrained extractor. They further proposed a method to protect noise-to-image GANs, which embeds the watermark in the generator during training [31]. Wu et al. [9] proposed a framework similar to [31] but for image-to-image models, where watermark embedding and model functionality are jointly trained. Alternatively, Zhang et al. [10] decoupled the two tasks by designing a dedicated hiding network that is encapsulated together with the original model. Our focus in this paper is box-free methods, and we will implement the state-of-the-art [9] and [10] for our experiments.
@@
Figure 1: Flowchart of box-free watermarking for image-to-image models, where image denoising is used as an example. It consists of an image processing network (IPNet), a hiding network (HNet), an extraction network (ENet), and a discriminator (D), respectively.
2.4 Watermark Removal Attacks
@@
The watermark removal attack generally aims to remove watermarks embedded in model parameters, behaviors (backdoor), or outputs, and thus result in failure of extraction or watermark verification. It can be either a white-box or a black-box attack, with or without the access to the victim model’s internal information. Lukas et al. [15] comprehensively investigated the robustness of model watermarks against various removal attacks including image preprocessing, white-box model modification, and black-box model stealing. In this spirit, Wang et al. [32] incorporated adversarial perturbations for preprocessing-based removal, while Guo et al. [16] proposed an improved fine-tuning-based white-box removal method. However, all these methods are only applicable to classification models.
@@
We also note that the robustness against potential removal attacks has been considered in existing watermarking works for generative models, but these attacks are merely common image preprocessing, e.g., cropping, blurring, compression, and adding noise [31], [9]. Recently, Liu et al. [33] developed a dedicated black-box remover against deep hiding models based on image inpainting. It first removes a few patches in a container image and then repairs the image via deep inpainting. However, this method requires the image before watermarking (equivalent to the output of IPNet) to supervise the inpainter, which is not available in real-world situations wherein the IPNet and HNet are encapsulated together–the intermediate output from IPNet is not observable and only the output of HNet (container) is observable. Therefore, our paper is the first work to address the vulnerability of box-free model watermarks against black-box removal attack in practical situations.
@@
TABLE I: The three scenarios (Scen.) considered in this paper, in which all models are in black box. “Activ.” stands for activation.
@@
3 Problem Formulation
3.1 Image-to-Image Model Box-Free Watermarking
To make this paper complete, we first present the fundamentals of box-free watermarking. The flowchart of the existing box-free watermarking is presented in Fig. 1 with image denoising as an example. It consists of an image processing network (IPNet), a hiding network (HNet), an extraction network (ENet), and a Discriminator (D), respectively. We assume that the original images $a_{1},a_{2},\ldots,a_{M}$ belong to domain $A$, and corresponding processed (denoised, in this example) images $b_{1},b_{2},\ldots b_{M}$ belong to domain $B$. The IPNet is thus trained to minimize the task loss, given by
$\mathcal{L}_{\text{Task}}=\left\|\text{IPNet}\left(a_{i}\right)-b_{i}\right\|^{2}_{2},$ (1)
where $1\leq i\leq M$. In order to safeguard the model, particularly against surrogate attacks utilizing surrogate models, HNet is employed to embed an invisible watermark, denoted by $\xi$ (IEEE mark image in Fig. 1 for example), into $b_{i}$. Specifically, $b_{i}$ and $\xi$ are concatenated along the channel dimension and fed into HNet, which then yields the watermarked $b^{\prime}_{i}$ in domain $B^{\prime}$. Meanwhile, ENet is supposed to generate the watermark image from $b^{\prime}_{i}$ as close to $\xi$ as possible and generate an all-white blank image, denoted by $\xi_{0}$, from non-watermarked images [^2] The predefined ENet output images for non-watermarked inputs can be noise, solid colors, or other low entropic images. In this paper, we use an all-white empty image as shown in Fig. 1.. To achieve this functionalities, HNet and ENet are then jointly trained with the joint loss given by
where
are the hiding loss, watermark loss, and clean loss, respectively, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are weighting hyperparameters, and $N$ is the number of pixels. There exist a few optional loss functions, e.g., the perceptual loss[34] and consistency loss [8] for some performance improvements, while $D$ can also be omitted. It has been show in [9] and [10] that ENet can extract the hidden watermark image not only from $b^{\prime}$s generated by HNet, but also from surrogate model outputs. This indicates that one can not only verify images generated by IPNet but also trace back surrogate models to approximate the functionality of IPNet.
Figure 2: Flowchart of the victim model and the proposed attacks. (a) and (c) are the EGG remover and transferable remover, respectively, while (b) is the victim model. Images processed by EGG RNet and Transferable RNet will be watermark-free, corresponding to all-white outputs from ENet. Note that the proxy models are private.
@@ 3.2 Threat Model
For the ease of discussion, we consider the victim model consisting of IPNet and HNet as a single entity called operation network (ONet). The three scenarios are then summarized in Table I, which defines our threat models. Combining Fig. 1 and Table I, we note that in all scenarios, the adversary can only observe the images to be processed in domain $A$ and the processed and watermarked images in domain $B^{\prime}$, while the intermediate results (processed but non-watermarked images in domain $B$) are strictly private.
In practice, Scenario 2 can be the case that ENet is provided as a black-box cloud service and can be used by end users to verify if an image is generated by IPNet, while Scenario 3 corresponds to the situation in which ENet is kept private by an authority for watermark verification.
For Scenarios 1 and 2, the attack designs a removal network (RNet) such that 1) if the input of RNet a watermarked image (output of ONet in domain $B^{\prime}$), then the output is the watermark removed image, and 2) if the input is a non-watermarked image, then the output does not contain the watermark, either. In both cases, the output image quality should be close to that of the input. For Scenario 3, since the observable information is minimal, we resort to transferability and design a proxy remover denoted by RNet, which is then used in the same way as RNet.
@@ 4 Proposed Attacks
The flowchart of the proposed attacks against the victim model is shown in Fig. 2, where IPNet is omitted for brevity. We note that querying HNet is required in the EGG attack but not in the transferable attack. The components that are used to perform the attack are the EGG RNet and Transferable RNet, respectively. For all scenarios, RNet takes an image $b_{i}^{\prime}$ from domain $B^{\prime}$, concatenated with the all-white image $\xi_{0}$, as input and attempts to generate the watermark-removed output $b_{i}^{\prime\prime}$ in domain $B^{\prime\prime}$. For a non-watermarked input, RNet will yield a high quality non-watermarked output, still in domain $B^{\prime\prime}$. We thus denote
@@
where $\oplus$ denotes channel-wise concatenation. In what follows, we first present the EGG removers under Scenarios 1 and 2, and then the transferability based remover.
@@ 4.1 Scenarios 1 and 2: EGG Removers
Since ONet performs two tasks (via IPNet and HNet) on a query image, its output $b_{i}^{\prime}$ is then the processed and watermarked image. An ideal removal attack should compensate HNet only, while keeping the image alteration due to IPNet intact. Unfortunately, the intermediate result $b_{i}$ is inaccessible, making it impossible to reverse the functionality of HNet by observing its input ($b_{i}$)-output ($b_{i}^{\prime}$) pairs. Under this stringent but practical attack condition, we are instead interested in exploring whether a self-supervised image reconstructor, supervised only by the observable $b_{i}^{\prime}$, can do the job. Therefore, in Scenarios 1 and 2, we minimize the following loss function when training RNet,
@@
$\beta_{1}$ and $\beta_{2}$ are weighting hyperparameters. $\mathcal{L}_{\text{Quality}}$ preserves the image quality after watermark removal, while $\mathcal{L}_{\text{Remove}}$ ensures that the output $b_{i}^{\prime\prime}$ does not contain the watermark. Note that (8) is practically obtainable since it only depends on the observable $b_{i}^{\prime}$ from ONet API. Unfortunately, the problem arises when evaluating (9) which involves querying ENet using the output of RNet, i.e., $b_{i}^{\prime\prime}$. Specifically, it requires the gradients of ENet to be back-propagated during the learning of RNet parameters. However, ENet is in a black-box, only its output is available.

### v3

@@ 1 Introduction
Common goals of DNN intellectual property protection are ownership verification and model stealing (also known as model extraction or surrogate attack) tracking. Both goals can be effectively achieved through DNN watermarking [5, 6, 7]. Existing DNN watermarking methods can be classified into white-box, black-box, and box-free methods [5]. In white-box methods, the watermark is embedded directly into the model’s parameters, which requires access to the model’s internal structure for watermark extraction. While this approach allows for accurate watermark extraction due to direct access to the model parameters, it sacrifices the model privacy. Meanwhile, providing access to a model’s structure and parameters for extraction may also violate security protocols, leading to the potential alteration or removal of the watermark. In contrast, black-box methods circumvent the need to access the model’s internal information. Instead, they employ specific inputs, referred to as triggers, which can alter the model’s behavior. This is also termed as backdoor watermarking, and the verification is done by analyzing model inputs and outputs.
@@
The existing white- and black-box methods are primarily employed for the protection of classification models. We note that classification models only generate low entropic outputs, e.g., the categorical confidence scores, or even merely the top-1 label. This aligns with the essence of white- and black-box methods that embed watermarks into model parameters and behaviors, respectively. In contrast, for generative models, e.g., image processing models we focus on in this paper, the outputs contain high entropic image content, allowing the models to generate watermarked outputs, known as box-free watermarking [8]. In this type of methods, a watermark hiding network and an extraction network are respectively designed and can be trained jointly with the original model. Then, the protected model (original model plus hiding network) generates processed and watermarked images, while the extraction network, when queried by (stolen) model-generated images, yields the watermark, or otherwise yields a noisy or non-informative pattern. Box-free methods excel at their flexibility as they do not require access to model parameters. The output images are also called containers [8].
@@
The existing works on box-free methods have focused on watermark fidelity, robustness against surrogate attacks, robustness against normal image processing, and extension to different generative tasks [9, 10, 11]. Despite the current research progress, in this paper, we make the first attempt to reveal an overlooked vulnerability of box-free methods against black-box watermark removal attacks. We focus exclusively on deep generative models for low-level image processing tasks and consider three scenarios. Scenario 1: The victim model, consisting of an image processing network (IPNet), a watermark hiding network (HNet), and a watermark extraction network (ENet), is secured in a black box, whose API are available for query, and ENet only uses ReLU as activation. We then propose an extractor-gradient-guided (EGG) watermark removal attack and show its effectiveness. Scenario 2: It reflects a more challenging yet practical situation which is almost the same as Scenario 1 except that ENet uses arbitrary and unknown activation functions. In this scenario, ENet’s gradients can be estimated via querying, according to studies in adversarial attacks [^1] In black-box adversarial attack literature [12, 13, 14], the gradient information of the black-box victim model is estimated via querying. While this is beyond the scope of this paper, it validates our Scenario 2 in which the gradients of ENet can be reverse-engineered., and we develop an EGG remover under this condition. Scenario 3: It is the most challenging scenario in which the victim model is in a black box while ENet API is not provided. Since the original functionality of IPNet is known, we resort to a transferability-based attack, which trains a private set of proxy models denoted by $\text{HNet}^{\prime}$, and $\text{ENet}^{\prime}$, respectively. Then, we show that the remover trained for the proxy models can be transferred to attack the victim model.
@@
In the above three scenarios, the proposed EGG removers, as well as the transferable remover, when fed with container (processed and watermarked) images, can effectively remove the embedded watermarks while preserving the image quality. We note that the proposed removers are also image-to-image models, which turn a container image into a watermark-free one. This differs from existing image-processing-based removal (e.g., adding noise, performing transformations, etc., usually less effective) [15] and white-box ones (e.g., fine-tuning and compression, less practical than black-box ones) [16]. The contributions of this paper are summarized as follow:
@@
  • We reveal the vulnerabilities of the existing box-free model watermarking methods against black-box removal attacks. It takes a distinctive perspective compared to most existing works dealing only with white-box removal such as fine-tuning and compression. This indicates that even black-box protection cannot effectively prevent watermark removal.
@@ 2.2 Black-Box Watermarking
Black-box model watermarking is also known as backdoor watermarking, whose watermark extraction process only requires the observation of the trigger-output relationships. Most existing backdoor watermarks are in the form of trigger-label mappings designed specially for deep classification models. The triggers can be both in- and out-of-distribution samples [18, 23, 24, 25, 26], while the trigger labels can be drawn from either the original categories or a novel category [27, 28]. However, Huang et al. [11] have pointed out the limitations of existing black-box methods in protecting generative models, noting that these methods are not applicable to noise-to-image models. In addition, black-box methods are unable to protect the generated images, which are also considered intellectual properties. In summary, while black-box watermarking could be more secure than white-box watermarking, it is unsuitable for safeguarding generative models. We note that Quan et al. [29] have made the first attempt to backdoor watermark generative models.
@@ 2.3 Box-Free Watermarking
Box-free watermarking was motivated by the need for protecting the high-entropic generated images [5]. The term “box-free” originates from the minimal dependency of watermark extraction on the model itself. In this line of research, Yu et al. [30] proposed an indirect approach to protect image-to-image generative adversarial networks (GANs). It first watermarks the training data and then, based on transferability, extracts the watermarks from images generated by the trained GANs using a pretrained extractor. They further proposed a method to protect noise-to-image GANs, which embeds the watermark in the generator during training [31]. Wu et al. [9] proposed a framework similar to [31] but for image-to-image models, where watermark embedding and model functionality are jointly trained. Alternatively, Zhang et al. [10] decoupled the two tasks by designing a dedicated hiding network that is encapsulated together with the original model. We note that the box-free watermarking considered in this paper differs from the subject of “hiding images within images” [32], as the former aims for model stealing tracking but the latter is for high-capacity data hiding. We will implement the state-of-the-art box-free methods [9] and [10] for experiments.
@@
Figure 1: Flowchart of box-free watermarking for image-to-image models, where image denoising is used as an example.
2.4 Watermark Removal
@@
Watermark removal attacks aim to remove watermarks embedded in model parameters, behaviors (backdoor), or outputs, and thus result in failure of extraction. It can be either a white-box or a black-box attack, with or without the victim model’s internal information. Lukas et al. [15] comprehensively investigated the robustness of model watermarks against removal attacks including image preprocessing, white-box model modification, and black-box model stealing. In this spirit, Wang et al. [33] incorporated adversarial perturbations for preprocessing-based removal, while Guo et al. [16] proposed an improved fine-tuning-based white-box removal attack. However, all these methods are only applicable to classification models and white-box setting.
@@
For image processing models, removal is usually related to robustness, and existing works have considered watermark robustness against normal image preprocessing, e.g., cropping, blurring, compression, and adding noise [31], [9]. Recently, Liu et al. [34] developed a model agnostic lattice attack and an inpainting-based black-box remover. The lattice attack replaces image pixels at specific locations by random values to remove the watermark, but it results in visible artifacts. The inpainting-based attack, on the other hand, first removes a few image patches and then repairs them via deep inpainting. However, it requires the image before watermarking, i.e., the output of IPNet, to supervise the inpainter, which is not available in our black-box setting.
Other types of watermark removal attacks are noted here for completeness. Jiang et al. [35] proposed the so-called WEvade-B-S, but with limited performance in their experiments, and WEvade-B-Q, a recursive attack needing multiple iterations to attack a single image, which is time and computational inefficient. You et al. [36] designed a dedicated remover to remove spread spectrum (SS) image watermarks, but it is a known message attack (KMA) requiring watermark information, which is not applicable in our black-box setting. Additionally, Niu et al. [37], Tian et al. [38, 39] studied the removal of visible watermarks, whereas our work targets the more challenging task of removing invisible watermarks. We also notice that the ambiguity attacks can confuse ownership verification by introducing multiple watermarks, but here we focus on the more formidable challenge of complete watermark removal.
TABLE I: Summary of symbols and abbreviations.
Symbol/Abbr. Meaning
IPNet Image processing network
HNet Hiding network
ONet Operation network (IPNet + HNet)
ENet Extraction network
D Discriminator
RNet Removal network
$a_{i}\in A$ To-be-processed image
$b_{i}\in B$ Processed unmarked image
$b_{i}^{\prime}\in B^{\prime}$ Processed marked image
$b_{i}^{\prime\prime}\in B^{\prime\prime}$ Output of RNet
$\xi$ Embedded watermark
$\xi_{0}$ Null watermark (all-white)
$\xi_{\text{Overwrite}}$ Watermark for overwriting
3 Problem Formulation
3.1 Image-to-Image Model Box-Free Watermarking
The flowchart of the existing box-free watermarking methods is presented in Fig. 1 with image denoising as an example, and the list of main symbols and abbreviations is provided in Table I. We assume that the original images $a_{1},a_{2},\ldots,a_{M}$ belong to domain $A$, and corresponding processed (denoised, in this example) images $b_{1},b_{2},\ldots b_{M}$ belong to domain $B$. To protect the pretrained IPNet from model stealing attacks, HNet is employed to embed an invisible watermark, denoted by $\xi$ (the IEEE mark image in Fig. 1 for example), into IPNet output $b_{i}$. Specifically, $b_{i}$ and $\xi$ are channel-wise concatenated and fed into HNet, which then yields the watermarked $b^{\prime}_{i}$ in domain $B^{\prime}$. Meanwhile, ENet is supposed to generate the watermark image from $b^{\prime}_{i}$ as close to $\xi$ as possible and generate an all-white image, denoted by $\xi_{0}$, from non-watermarked images [^2] The predefined ENet output images for non-watermarked inputs can be noise, solid colors, or other low entropic images. In this paper, we use an all-white image as shown in Fig. 1.. To achieve these functionalities, with the pretrained IPNet frozen, HNet and ENet are jointly trained by minimizing
where
are the hiding loss, watermark loss, and clean loss, respectively, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are weighting hyperparameters, and $N$ is the number of pixels. There exist a few optional loss functions, e.g., the perceptual loss[40] and consistency loss [8] for some performance improvements, while $D$ can also be omitted. It has been show in [9] and [10] that ENet can extract the hidden watermark image not only from $b^{\prime}$s generated by HNet, but also from surrogate model outputs. This indicates that one can not only verify images generated by IPNet but also trace back surrogate models the approximate IPNet.
Figure 2: Flowchart of the victim model and the proposed attack, where ONet and ENet are secured in black-box, with only their API available to attacker. Images processed by RNet will be watermark-free, corresponding to all-white outputs from ENet.
@@ 3.1 Image-to-Image Model Box-Free Watermarking
TABLE II: The three scenarios (Scen.) considered in this paper, in which all models are in black box. “Activ.” stands for activation.
@@ 3.2 Threat Model
For the ease of discussion, we consider the victim model consisting of IPNet and HNet as a single entity called operation network (ONet). The three scenarios are then summarized in Table II that defines our threat model. Combining Fig. 1 and Table II, we note that in all scenarios, the attacker can only observe the to-be-processed images in domain $A$ and the processed and marked images in domain $B^{\prime}$, while the intermediate results ($b_{i}\in B$) are strictly private. The ReLU activation condition in Scenario 1 may not hold in practice, but this scenario is considered for theoretical analysis and is followed up by the more practical Scenario 2. In practice, Scenario 2 can be the case that ENet is provided as a black-box cloud service and can be used by end users to verify if an image is generated by IPNet, while Scenario 3 corresponds to the situation in which ENet is kept private by an authority for watermark verification.
For Scenarios 1 and 2, the attacker designs the RNet such that 1) if RNet input is a watermarked image, then the output is the watermark-removed image, and 2) if RNet input is a non-watermarked image, then the output does not contain the watermark, either. In both cases, the output image quality should be close to that of the input. For Scenario 3, since the observable information is minimal, we resort to transferability and design a proxy remover denoted by RNet, which is then used in the same way as RNet.
Figure 3: Flowchart of the training process for the proposed (a) EGG RNet and (b) Transferable RNet. Note that the proxy models are private and are frozen during the training of the Transferable RNet.
@@ 4 Proposed Attacks
The general flowchart of the proposed attack against the victim model is shown in Fig. 2. while the flowcharts of the training process for EGG RNet and Transferable RNet are presented in Fig. 3. IPNet in Fig. 3 is omitted for brevity. We note that querying HNet is required in the EGG attack but not in the transferable attack. The components that are used to perform the attack are the EGG RNet and Transferable RNet, respectively. For all scenarios, RNet takes an image $b_{i}^{\prime}$ from domain $B^{\prime}$, concatenated with the all-white image $\xi_{0}$ as input, and attempts to generate the watermark-removed output $b_{i}^{\prime\prime}$. For a non-watermarked input, RNet will yield a high quality non-watermarked output. This is modeled by
@@
where $\oplus$ denotes channel-wise concatenation. In what follows, we first present the EGG removers under Scenarios 1 and 2, and then the transferable remover under Scenario 3.
@@ 4.1 Scenarios 1 and 2: EGG Removers
Since ONet performs two tasks via IPNet and HNet on a query image, its output $b_{i}^{\prime}$ is then the processed and watermarked image. An ideal removal attack should compensate HNet only, while keeping the image alteration due to IPNet intact. Unfortunately, the intermediate result $b_{i}$ is inaccessible, making it impossible to reverse the functionality of HNet by observing its input ($b_{i}$)-output ($b_{i}^{\prime}$) pairs. Under this stringent but practical attack condition, we are instead interested in exploring whether a self-supervised image reconstructor, supervised only by the observable $b_{i}^{\prime}$, can do the job. Therefore, in Scenarios 1 and 2, we minimize the following loss function when training RNet,
@@
$\beta_{1}$ and $\beta_{2}$ are weighting hyperparameters. $\mathcal{L}_{\text{Quality}}$ preserves the image quality after watermark removal, while $\mathcal{L}_{\text{Remove}}$ ensures that the output $b_{i}^{\prime\prime}$ does not contain the watermark. Note that (7) is practically obtainable since it only depends on the observable $b_{i}^{\prime}$ from ONet API. Unfortunately, the problem arises when evaluating (8) which involves querying ENet. Specifically, it requires the gradients of ENet to be back-propagated during training of RNet. However, ENet is in a black-box, and only its output is available.